What is Sn of the geometric series with a1=4, an=256, and n=4?
A)84
B) 336
C) 340
D)260

Answer by richwmiller(17219)

First we need to find the ratio
Tn = t * r^(n - 1)
256 = 4 * r^(3)
r=4
S=t*(1 - r^n)/(1 - r)
S=4*(1 - 4^4)/(1 - 4)=340
C) 340